OpenStudy (anonymous):

Can someone help me solve these questions? a) If (x-2) is a factor of 2x3-3x+k, find k. b) When ax3-4x2+5x-3 is divided by (x+2) and (x-1), the remainders are equal. Find a.

OpenStudy (anonymous):

if x-2 is a factor of 2x3-3x+k, then 2(2)^3 -3(2) +k =0

OpenStudy (anonymous):

use if F(x) is divided by x-a then the Remainder is given by F(a) in case if x-a is a factor of F(x) then F(a) =0

OpenStudy (anonymous):

Can you please help me solve it mat?

hero (hero):

@rcatabijan, Did you find the value of k yet?

OpenStudy (anonymous):

Not yet, I need help in doing so..

OpenStudy (anonymous):

is it because you moved "k"? to the other side?

OpenStudy (anonymous):

by shifting it to the other side right? would the answer be k = -2(2)^3 + 3^2? we can say we found k?

OpenStudy (anonymous):

can you please help me simplify it

OpenStudy (anonymous):

k = 16 + 9 k = 25

OpenStudy (anonymous):

yeah?

OpenStudy (anonymous):

oh yes because of the negative.. so it would be -7

OpenStudy (anonymous):

how about for b), how would I go about starting the question?

OpenStudy (mathmale):

Mind staying with (a) for a few minutes longer? Have you checked to determine whether x-2 is truly a factor of 2x^3-3x+k when k=-7?

OpenStudy (mathmale):

But, my Hero, wouldn't -3(2)=-6, not -9?

hero (hero):

Ugh, stems from a mistake I made earlier.

hero (hero):

k = -2(2)^3 + 3(2) k = -2(8) + 6 k = -16 + 6 k = -10

OpenStudy (mathmale):

I found that synthetic div. works wonderfully well in solving this problem. Having solved it, I noticed that the previous posters came up with k=-7, which is incorrect. k=-10 is right.

hero (hero):

What's wrong with @matricked method?

hero (hero):

By the way, you never got back to me regarding the file I sent you earlier I notice.

OpenStudy (mathmale):

Have you read part B, and if so, have you begun solving it? I'm curious how you'd proceed. Again I'm starting out with synth. div. because this method is usually so short and efficient. I did look at that file and will look at it again. I believe I found one small error similar to your -3(2)=-9.

hero (hero):

I know. I wrote that mistake while trying to avoid it. I know exactly what you're talking about.

hero (hero):

@mathmale, if you keep referring back to that mistake so much, I'm may feel the urge to delete your comment. Constantly referring back to that mistake isn't helping anyone as much as it is insulting.

hero (hero):

You acknowledged the mistake, now move on.

OpenStudy (mathmale):

Why not correct the doc and then share it with me again? Another thought I had while reading it was that some might be more willing to read it were you to provide some verbal explanations and directions instead of just the math symbolism. Frankly, I'm taken aback at your outburst of anger. I had had NO intention of dwelling on it. I'd written the above suggestion when your angry message came through. Perhaps this is not the right venue to be talking about your method.

hero (hero):

I never said I was angry about anything. I'll get back to you later though.

OpenStudy (mathmale):

(Somewhat at a loss for words.) All right.